Circular No. 6317

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996C IN MCG +8-25-47 J. Mueller reports her discovery of a supernova (mag ∼ 18) on a IV-N

plate taken on Feb. 15 UT by K. M. Rykoski and herself with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. SN 1996C is located 14′′ north of the center of MCG +08-25-047 (α = 13^{h}50^{m}.8, δ = +49^{o}20′; equinox 2000.0). No object appears at this position on original Palomar Sky Survey prints or on second Sky Survey transparencies.

P. Garnavich, A. Riess, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that an image obtained by J. Luu with the Mt. Hopkins 1.2-m telescope on Feb. 17.5 UT shows the supernova at V ∼ 16. At about this same time, D. Koranyi obtained a spectrogram with the 1.5m Tillinghast telescope, showing SN 1996C to be a type-Ia supernova near maximum. A weak Hα emission line superimposed on the spectrum has a redshift of 0.027.

SUPERNOVA 1996D IN NGC 1614 L. Drissen, C. Robert, Y. Dutil, and J.-R. Roy, Département de Phy-

sique, Université Laval, report the discovery of a supernova on an image and spectrogram obtained at the Canada-France-Hawaii telescope on Feb. 9. A preliminary inspection of the partially-calibrated data by A. V. Filippenko (University of California at Berkeley) suggested that this might be a type-Ic supernova. SN 1996D is located in a spiral arm, ∼ 6′′.6 east of the bright nucleus of NGC 1614 (α = 4^{h}34^{m}.0, δ = −8^{o}35′, equinox 2000.0). This new source is not visible in previous images, the latest known being a Hubble Space Telescope WFPC2 image (F606W) collected on 1994 Dec. 11.

E. Cappellaro, Osservatorio Astronomico di Padova; and F. Patat, European Southern Observatory (ESO) report: “On Feb. 18.0 UT, we observed SN 1996D with the 2.2-m ESO telescope (+ EFOSC2). Based on a low-resolution (3.5 nm) spectrum (range 380–900 nm), we confirm that this is indeed a type-Ic supernova near maximum light. The most prominent features are O I at 777.3 nm and the Ca II infrared triplet, both showing a P-Cyg profile and indicating an expansion velocity of ∼ 8000 km/s. The recession velocity derived from the narrow Hα and S II emissions of the underlying H II region is 4750 km/s. Preliminary photometry gives V = 18.2, V − R = +0.7.”

1996 February 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green